OpenStudy (anonymous):

how does one compute the following: (d^2y/dt^2)-t-y=0 Please use the Huen method (modified Eulers)

OpenStudy (anonymous):

my question is how do I advance the first order derivative?

OpenStudy (anonymous):

start with \[w(y,t)={dy\over dt}\\ {dw\over dt}-t-y(t)=0\] this form can now be implemented using Euler's or modified Euler's or any other linear approximation methods.

OpenStudy (anonymous):

I agree thank you. how to I advance w?

OpenStudy (anonymous):

treat the new differential as a separate problem between two iterations of "y"

OpenStudy (anonymous):

so, w advances, then y advances, then w again.. and then y again.. and so on

OpenStudy (anonymous):

Ah, so yo then wo then y1 and w1 Thank you

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

think of "y" as position then \(w={dy\over dt}\)is your velocity and \({dw\over dt}={d^2y\over dt^2}\)is your acceleration function.

OpenStudy (anonymous):

notice that y -> function of "t" and w-> function of "t" and "y"

OpenStudy (anonymous):

@garrett_payne now, how about being courteous and award?

OpenStudy (anonymous):

lol, just joined. How do I do that?

OpenStudy (anonymous):

jsut say best response :)

OpenStudy (anonymous):

issue; what happens when dy/dt is 0 the problem is as follows \[\frac{ d ^{2} y}{ dt ^{2} }-t+y = 0 \] I haved called \[\frac{dy }{ dt }=z \] and step size is 0.1 y(0)=2 z(0)=0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

could you show 1 iteration?

OpenStudy (anonymous):

ok \[ h=\text{step size}\\ \frac{dz}{dt}=t-y=f_1(t,y)\\ \frac{dy}{dt}=z(t,y) \] k^th Iteration steps: \[ t_{k+1}=k_k+h\\ y_{k+1}=y_k+hz_k\qquad\qquad\text{(first approx.}\\ z_{k+1}=z_k+\frac{h}{2}\left[f_1(t_k,y_k)+f_1(t_{k+1},y_k)\right]\\ y_{k+1}=y_k+\frac{h}{2}\left[z_k+z_{k+1}\right]\qquad\qquad\text{(second approx.} \]

OpenStudy (anonymous):

correction... in the second term of \(z_{k+1}\), I wrote \(y_k\) by mistake instead of \(y_{k+1}\)

OpenStudy (anonymous):

I caught that, Thank you so much My numerical operations prof is an retrice

OpenStudy (anonymous):

for accuracy, you may choose to perform the last two steps twice to improve accuracy

OpenStudy (anonymous):

do you mean, replugging the obtained value in the same equation?

OpenStudy (anonymous):

yes. coz, we approximated the y_{k+1} the first time...